OpenStudy (zepp):

Post the whole process to solve this please :D \[\large \int_{-1}^{0} \frac{1}{2}x+\frac{1}{2}dx\] Thanks!

OpenStudy (anonymous):

\[\int\limits \frac{1}{2}xdx+\frac{1}{2} \int\limits dx\] \[(\frac{1}{4}x^2+\frac{1}{2}x)|^0_{-1}\] \[[(\frac{0^2}{4}+\frac{0}{2})-(-\frac{1}{2})]=\frac{1}{2}\]

OpenStudy (zepp):

I got \(\large \frac{1}{4}\) using geometry.. :|

OpenStudy (anonymous):

That's because I left out part of the last integral reult in the last line, ha \[[(\frac{0^2}{4}+\frac{0}{2})-(-\frac{1}{4}(-1)^2-\frac{1}{2})]=\frac{1}{4}\]

OpenStudy (anonymous):